# Extended symmetry analysis of (1+2)-dimensional fine Kolmogorov backward equation

### v1

@@ 7 Generalized reductions
[…]
@@ Appendix A Pseudogroups
We first present an (enhanced) definition of pseudogroup of transformations of a topological space which is often used in the literature [28, Section 1.1] and can be further specified for topological spaces with additional structures (i.e., localé) such as manifolds, and then modify it. A pseudogroup of transformations on a topological space $X$ is a set $G$ of homeomorphisms between open subsets of $X$ with the operation of map composition that satisfy the following properties:
@@
An (algebraic) groupoid is a nonempty set $G$ together with a unary operation (inverse) :${}^{-1}\colon G\to G$ and a partial binary operation111***” is not a usual (total) binary operation since it is not necessarily defined for all pairs of elements of $G$. (multiplication) *$*\colon G\times G\rightharpoonup G$ satisfying the following properties:
@@
([16, p. 6]) An inverse semigroup is a nonempty set $S$ with an associative binary operation *$*\colon S\times S\to S$ such that for every $a\in S$, there exists a unique $a^{-1}\in S$, called the inverse of $a$, satisfying $a^{-1}*a*a^{-1}=a^{-1}$ and $a*a^{-1}*a=a$. An inverse monoid $M$ is an inverse semigroup with neutral element $e\in M$, $e*a=a*e=a$ for all $a\in M$.
@@
An element $a$ of a set with a binary operation *** is called an idempotent if $a*a=a$. One can completely characterize the set of idempotents in any pseudogroup.
@@ Appendix B Two-by-two matrix groups as semidirect products
where $\iota$ is the inclusion homomorphism. We need to show that the above sequence splits. In other words, there exists a right section of $\det$, $\nu\colon{\rm U}(\mathbb{Z})\to{\rm SL}^{\pm}(2,\mathbb{R})$, such that ${\rm SL}^{\pm}(2,\mathbb{R})={\rm SL}(2,\mathbb{R})\nu({\rm U}(\mathbb{Z}))$ and ${\det}\circ\nu={\rm id}_{\rm U(\mathbb{Z})}$. To satisfy the first condition, we should require that $\det\nu(-1)=-1$, and since $\nu(-1)^{2}={\rm diag}(1,1)$, the matrix $\nu(-1)$ is diagonalizable with the eigenvalues ±$\pm 1$, so the Jordan normal form for $\nu(-1)$ is ${\rm diag}(1,-1)$. Taking the map $\nu\colon{\rm U}(\mathbb{Z})\to\rm{SL}^{\pm}(2,\mathbb{R})$ defined on the generator of ${\rm U}(\mathbb{Z})$ by $-1\mapsto{\rm diag}(1,-1)$ as the canonical homomorphism, we get that ${\rm SL}^{\pm}(2,\mathbb{R})$ is the semidirect product of its subgroups ${\rm SL}(2,\mathbb{R})$ and $\nu({\rm U}(\mathbb{Z}))$, ${\rm SL}^{\pm}(2,\mathbb{R})={\rm SL}(2,\mathbb{R})\rtimes\nu({\rm U}(\mathbb{Z}))$. The conjugation in ${\rm SL}(2,\mathbb{R})$ by the matrix $\nu(-1)$ gives the desired involution $\varphi(\bar{1})$. ∎

### v2

@@ 7 Generalized reductions
The Lie-symmetry vector field $-\mathcal{P}^{t}-\mu\mathcal{I}$ belongs to the center $\mathfrak{z}$ of the algebra $\mathfrak{g}^{\rm ess}$ and thus the associated Lie-symmetry operator $\hat{\mathrm{P}}^{t}-\mu$ commutes with all elements of $\Upsilon_{\mathfrak{f}}$. This is why the solutions of (2) that are invariant with respect to $(({\rm D}_{t}-\mu)^{n}u)\partial_{u}$ cannot be related to generating solutions with iterative action of Lie-symmetry operators, cf. paragraph 1.1μ in Section 6.
@@ Appendix A Pseudogroups
We first present an (enhanced) definition of pseudogroup of transformations of a topological space which is often used in the literature [28, Section 1.1] and can be further specified for topological spaces with additional structures (i.e., localé) such as manifolds, and then modify it. A pseudogroup of transformations on a topological space $X$ is a set $G$ of homeomorphisms between nonempty open subsets of $X$ with the operation of map composition that satisfy the following properties:
@@
An (algebraic) groupoid is a nonempty set $G$ together with a unary operation (inverse) :${}^{-1}\colon G\to G$ and a partial binary operation111*” is not a usual (total) binary operation since it is not necessarily defined for all pairs of elements of $G$. (multiplication) $*\colon G\times G\rightharpoonup G$ satisfying the following properties:
@@
([16, p. 6]) An inverse semigroup is a nonempty set $S$ with an associative binary operation $*\colon S\times S\to S$ such that for every $a\in S$, there exists a unique $a^{-1}\in S$, called the inverse of $a$, satisfying $a^{-1}*a*a^{-1}=a^{-1}$ and $a*a^{-1}*a=a$. An inverse monoid $M$ is an inverse semigroup with neutral element $e\in M$, $e*a=a*e=a$ for all $a\in M$.
@@
An element $a$ of a set with a binary operation * is called an idempotent if $a*a=a$. One can completely characterize the set of idempotents in any pseudogroup.
@@ Appendix B Two-by-two matrix groups as semidirect products
where $\iota$ is the inclusion homomorphism.
We need to show that the above sequence splits. In other words, there exists a right section of $\det$, $\nu\colon{\rm U}(\mathbb{Z})\to{\rm SL}^{\pm}(2,\mathbb{R})$, such that ${\rm SL}^{\pm}(2,\mathbb{R})={\rm SL}(2,\mathbb{R})\nu({\rm U}(\mathbb{Z}))$ and ${\det}\circ\nu={\rm id}_{\rm U(\mathbb{Z})}$. To satisfy the first condition, we should require that $\det\nu(-1)=-1$, and since $\nu(-1)^{2}={\rm diag}(1,1)$, the matrix $\nu(-1)$ is diagonalizable with the eigenvalues ±$\pm 1$, so the Jordan normal form for $\nu(-1)$ is ${\rm diag}(1,-1)$.
Taking the map $\nu\colon{\rm U}(\mathbb{Z})\to\rm{SL}^{\pm}(2,\mathbb{R})$ defined on the generator of ${\rm U}(\mathbb{Z})$ by $-1\mapsto{\rm diag}(1,-1)$ as the canonical homomorphism, we get that ${\rm SL}^{\pm}(2,\mathbb{R})$ is the semidirect product of its subgroups ${\rm SL}(2,\mathbb{R})$ and $\nu({\rm U}(\mathbb{Z}))$, ${\rm SL}^{\pm}(2,\mathbb{R})={\rm SL}(2,\mathbb{R})\rtimes\nu({\rm U}(\mathbb{Z}))$. The conjugation in ${\rm SL}(2,\mathbb{R})$ by the matrix $\nu(-1)$ gives the desired involution $\varphi(\bar{1})$. ∎